# Ailing Yu1  Lan Yao2  Jingnan Liu1  Zhe Chen1  Jiajun Yin1 Yuan Wang1  Xinhao Liao1  Zhiling Ye1  Ji Li1  Yun Yue1 Hansong Xiao1  Hualei Zhou1  Chunxiao Guo1  Peng Wei1  Jinjie Gu1

### v2

@@
Figure 1: Overall performance of MedResearcher-R1 across three benchmarks. On MedBrowseComp, our MedResearcher-R1-32B achieves state-of-the-art 27.5% accuracy, surpassing o3-deepresearch (25.5%), Gemini-2.5-Pro-deepresearch (25.0%), and significantly outperforming search-only approaches (o3-search: 19.0%, Gemini-2.5-Pro-search: 14.0%). On general deep research tasks, we achieve competitive results on GAIA (53.4 vs. WebSailor-32B’s 53.2) and xBench (54.0 vs. WebSailor-32B’s 53.3).
@@ 1 Introduction
Our experimental results validate this approach. The trained model, MedResearcher-R1, achieves $27.5\%$ accuracy on MedBrowseComp, establishing a new state-of-the-art and substantially outperforming both the Qwen2.5-32B baseline and existing deep research systems. Notably, our medical specialization does not compromise general capabilities: on general agent benchmarks (GAIA: 53.4, xBench: 54), MedResearcher-R1maintains competitive performance comparable to GPT-4o-mini. Please refer to Figure 1 for an overview.
@@ 4.1 Cold Start with Supervised Fine-Tuning
Optimization proceeds using the AdamW optimizer (learning rate $\lambda=0.01$), with a cosine annealing schedule ($\eta_{\max}=3\times 10^{-7}$), for 3 epochs on 8 $\times$ H200 GPUs. This ensures rapid exploration of diverse trajectories and convergence to well-calibrated agent policies.
@@ 5.1 Benchmarks
  • MedBrowseComp (Chen et al., 2025b) is a recently proposed benchmark specifically designed to evaluate LLM agents’ capabilities in retrieving and synthesizing medical evidence from multiple web sources. This benchmark presents agents with open-ended clinical questions that necessitate multi-step reasoning, strategic information gathering, and effective utilization of web browsing APIs to construct comprehensive medical assessments.
@@
  • XBench-DeepSearch (Chen et al., 2025a) is an extensive multi-domain agent evaluation suite that systematically assesses tool-use capabilities across diverse open-domain tasks. The benchmark encompasses a broad spectrum of scenarios, including fact-checking, comparative analysis, web browsing-based reasoning, and complex information synthesis tasks, providing a comprehensive evaluation of an agent’s ability to navigate and utilize various tools in real-world problem-solving contexts.
@@ 5.2 Main Results
As shown in Table 1, our tool-augmented agent achieves new state-of-the-art performance on the MedBrowseComp benchmark, achieving a pass@1 accuracy of $27.5\%$ and outperforming both previous best agents and the Qwen2.5-32B baseline. The supervised fine-tuning (SFT) stage already delivers notable gains, while subsequent reinforcement learning further improves decision quality and tool orchestration efficiency.
@@
Model o3 search gemini2.5pro deepsearch o3 deepresearch claude-cua MedASearch-32B
@@ 5.3 Qualitative Analysis
Figure 5: Case study demonstrating the search-verify-synthesize paradigm: The agent performs multiple verification rounds across medical sources, ensuring information consistency before synthesis. Baseline agents (shown in gray) terminate prematurely after initial search, while our approach (blue) continues until achieving high confidence through cross-validation.

### v3

# Ailing Yu1  Lan Yao2  Jingnan Liu1  Zhe Chen1  Jiajun Yin1 Yuan Wang1  Xinhao Liao1  Zhiling Ye1  Ji Li1  Yun Yue1 Hansong Xiao1  Hualei Zhou1  Chunxiao Guo1  Peng Wei1  Junwei Liu1  Jinjie Gu1

@@
Figure 1: Overall performance of MedResearcher-R1 across three benchmarks. On MedBrowseComp, our MedResearcher-R1-32B achieves state-of-the-art performance with 27.5/50 correct answers, surpassing o3-deepresearch (25.5/50), Gemini-2.5-Pro-deepresearch (25.0/50), and significantly outperforming search-only approaches (o3-search: 19.0/50, Gemini-2.5-Pro-search: 14.0/50). On general deep research tasks, we achieve competitive results on GAIA (53.4 vs. WebSailor-32B’s 53.2) and xBench (54.0 vs. WebSailor-32B’s 53.3).
@@ 1 Introduction
Our experimental results validate this approach. The trained model, MedResearcher-R1, achieves a score of 27.5/50 on MedBrowseComp, establishing a new state-of-the-art and substantially outperforming both the Qwen2.5-32B baseline and existing deep research systems. Notably, our medical specialization does not compromise general capabilities: on general agent benchmarks (GAIA: 53.4, xBench: 54), MedResearcher-R1maintains competitive performance comparable to OpenAI o4-mini. Please refer to Figure 1 for an overview.
@@ 4.1 Cold Start with Supervised Fine-Tuning
Optimization proceeds using the AdamW optimizer (learning rate $\lambda=0.01$), with a cosine annealing schedule ($\eta_{\max}=3\times 10^{-7}$), for 3 epochs on 8 $\times$ H800 GPUs. This ensures rapid exploration of diverse trajectories and convergence to well-calibrated agent policies.
@@ 5.1 Benchmarks
  • MedBrowseComp (Chen et al., 2025b) is a recently proposed benchmark specifically designed to evaluate the capabilities of LLM-based agents in retrieving and synthesizing medical evidence from multiple web sources. This benchmark presents agents with open-ended clinical questions that necessitate multi-step reasoning, strategic information gathering, and effective utilization of web browsing APIs to construct comprehensive medical assessments.
@@
  • XBench-DeepSearch (Chen et al., 2025a) is an extensive multi-domain agent evaluation suite that systematically assesses tool-use capabilities across diverse open-domain tasks. The benchmark encompasses a broad spectrum of scenarios, including fact-checking, comparative analysis, web browsing-based reasoning, and complex information synthesis tasks, providing a comprehensive evaluation of the ability of LLM-based agents to navigate and utilize various tools in real-world problem-solving contexts.
@@ 5.2 Main Results
As shown in Table 1, our tool-augmented agent achieves new state-of-the-art performance on the MedBrowseComp benchmark, achieving a pass@1 score of 27.5/50 and outperforming both previous best agents and the Qwen2.5-32B baseline. The supervised fine-tuning (SFT) stage already delivers notable gains, while subsequent reinforcement learning further improves decision quality and tool orchestration efficiency.
@@
Model o3 search gemini2.5pro deepsearch o3 deepresearch claude-cua MedResearcher-R1-32B
@@ 5.3 Qualitative Analysis
Figure 5: Case study demonstrating the search-verify-synthesize paradigm: The agent performs multiple verification rounds across information sources, ensuring information consistency before synthesis. Baseline agents (shown in gray) terminate prematurely after initial search, while our approach (blue) continues until achieving high confidence through cross-validation.